OpenStudy (anonymous):

What is the simplified form o the fallowing expression 2x^2Y + 3x^2 + 4y + 3x^2y + 2y

OpenStudy (adamaero):

Hint: (2x^2Y + 3x^2y )+ 2y + 3x^2 + 4y

OpenStudy (anonymous):

You moved the other 2 whys from the 4y to the back to make a 4y i the back ? Don't understand

OpenStudy (anonymous):

@adamaero

OpenStudy (adamaero):

4y can be anywhere

OpenStudy (anonymous):

So now it is: 2x^2Y + 3x^2 + 2y + 3x^2y + 4y Can i mover ^ thoes in ^there ?

OpenStudy (adamaero):

\[6x ^{2y} + 3x ^{2} + 6y\] and there is one or two more simplifications

OpenStudy (anonymous):

How did u write it like that?

OpenStudy (adamaero):

lower left "Equation"

OpenStudy (anonymous):

oh cool!

OpenStudy (anonymous):

um you move the 3x^2 into the 6x ?

OpenStudy (anonymous):

@adamaero is that correct ?

OpenStudy (anonymous):

\[9x ^{4y} + 3x ^{2} + 6y ? That the finished one ?\]

OpenStudy (adamaero):

nope

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

is that completely wrong or just needs more simplifying ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

And do you know properties welll /?

OpenStudy (anonymous):

Well*

OpenStudy (adamaero):

Ok this is each step broken down: \[2x^{2y} + 3x ^{2y} = (2+3)x^{2y}\] (I accidentally multiplied above before) + 2y + 4y = (2+4)y + \[3x^2\] Then take a three out of that whole equation

OpenStudy (anonymous):

What do you mean take a 3 out of the equation. Do you mean the 3 from 3x^2 ?

OpenStudy (anonymous):

@adamaero

OpenStudy (adamaero):

Sorry I was going off my first mistake. The final answer is probably where I broke it down--no factoring out a three. I don't see it, but there might be completing the square left.

OpenStudy (anonymous):

so 2x^2y + 3x^2y = (2 + 3)x^2y ?

OpenStudy (adamaero):

ya

OpenStudy (anonymous):

oh okay Thank you! :) 

OpenStudy (anonymous):

lol